Calculate the price elasticity of demand using the mid-point Method from the following 

information.

Price 20 30 40

Quantity 90 60 45

Would your result be different if instead you consider the total revenue approach to 

calculate the price elasticity of demand? (5.75)

b. Show that a minimum wage set above the market clearing wage rate causes excess supply 

of labour. Why would a government impose minimum wage despite knowing its effect on 

unemployment? (6)

c. What is effect on producer surplus if taste changes in favor of organic food. Explain it 

diagrammatically. Will it affect producer surplus in the market for organic food?


1
Expert's answer
2021-03-31T07:21:57-0400

a) Let's estimate the point price elasticity


"E_{d} = \frac{\frac{(60-90)}{\frac{(60+90)}{2}}}{\frac{(30-20)}{\frac{(30+20)}{2}}}""= \frac{\frac{-30}{75}}{\frac{10}{25}} =" "- 1"

Total revenue approach:

"TR = P\times Q"

"TR = 1800; 1800; 1800"

If TR remains the same then "Mod(E_{d}) = 1"

Which means that both approaches have the same result

b)


The government would impose a minimum wage to keep up with inflation

C) since the change in taste will shift the demand downward producer surplus will fall. However producer surplus on a organic food market will rise due to higher demand